# CU-JADE: A Method for Traversing Extinction Jumps along the Line-of-Sight

### v1

@@ 3.2 Detection of Weak Jumps in D-A Plot
Considering the end of the iterative procedure in Steps 3 and 4 (as an example of Figure 1f), for a set of $D$-$A$ data with a single jump point, we can obtain $S_{\rm diff}^{0}$ according to the definition
@@
$\xi\left(\Delta A_{\lambda},\sigma\right)$ exhibits a complex analytic form. However, $\xi\left(\Delta A_{\lambda},\sigma\right)\propto\sigma$ when $\sigma\gtrsim\Delta A_{\lambda}$. We focus on analyzing two primary cases. When $\Delta A_{\lambda}\gg\sigma$, we define these features as significant extinction jump points. The algorithm facilitate efficient traversal of all jump points on the $D$-$A$ diagram (see examples in Figure 2). For the other case, due to intrinsic dispersion and uncertainty, the extinction of stars in Gaia data leads to a magnitude uncertainty of $\sigma$ that is comparable to most of the $\Delta A_{\lambda}$ values we can detect as jump points. In observations, these jump points are collectively referred to as “weak jumps”, and statistically, the number of these weak jumps accounts for the majority. Additionally, the $D$-$A$ diagram includes gentle slopes ($k$) that may be caused by diffuse and extended gas along the LoS, which exhibit more linear properties in $D$-$A$ space and present challenges for the CU-JADE method (see Appendix A.3). In the following subsections, we will address the influence of these factors on the results and optimize the parameter settings to identify these extinction features with weak jumps.
@@ 3.2.1 Comparable $\Delta A_{\lambda}$ and $\sigma$ for Weak Jumps
By Equation (6), (A7) and Figure A1, when $\Delta A_{\lambda}$ and $\sigma$ are the same magnitude, $\sigma$ has a major influence on the $S_{\rm diff}$. Approximately, we have $S_{\rm diff}~{}\propto~{}\sqrt{n}~{}\sigma,~{}\sigma\gtrsim\Delta A_{\lambda}$. The influence of $\Delta A_{\lambda}$ on $S_{\rm diff}$ is negligible compared to $S_{\rm diff}^{0}$. Combine Equation (5), we have
@@
The $CL$ for an extinction jump is proportional to the jump value and the square root of number density, but inversely proportional to $A_{\lambda}$ dispersion. Therefore, a larger sample size tends to make it easier to detect a jump point, as well as the larger $\Delta A_{\lambda}$. A large dispersion would smear out the jump when it is comparable to the $\Delta A_{\lambda}$ values. Since $\Delta A_{\lambda}$ and $\sigma$ depend upon the LoS extinctions and the uncertainties of data, the value of $CL$ now is only influenced by number density of samples in the calculation.
@@ 3.2.2 Two Modes for Stellar Sample Density Variation
The second mode (Mode II) adjusts the $CL$ threshold with the number density, scaling with $\sqrt{n}$, to mitigate the influence of star density. In Mode II, we refine our method to apply varying $CL$ thresholds in each split iteration, enabling the jump point detection in intervals with sparser star distribution. In cases with prior knowledge of the target area, we recommend the Mode II. It is important to control bias stemming from sample density. The number density $n$ is calculated based on the star samples number per unit length in a specific region and distance interval, $n=\frac{N}{d_{max}-d_{min}}$, where $N$ is the number of stars, $d_{max}$ and $d_{min}$ are the maximum and minimum distances of selected stars in each interval.
As discussed in Section 3.1, to enhance the $CL$ while mitigating false detections, we multiply $S_{\rm diff}$ by a factor $C$ ($C\sqrt{\beta n}$ for Mode II, $\beta$ is a constant for fine-tuning). $C$ is usually set to be $1-3$, considering $\Delta A_{\lambda}$ ranges from $\left[0.5,5\right]$ mag and $n$ ranges from $\left[0.1,10\right]~{}{\rm pc}^{-1}$. In the cases of large dispersion $\sigma$, or too small $\Delta A_{\lambda}$ due to weak extinctions, we have to do coarse tuning to the criteria by reducing the value of $C$, as $CL$ decreases rapidly during the iterations. We find that $C=1$ is more effective for weak jumps (see details in Section 4.1). Fine-tuning of the threshold can also be achieved by altering the percentile value of $S_{\rm diff}^{0}$ in shuffled $S_{\rm diff}$ (generally taken as 0.95). Mode II has to be complemented with both a lower and an upper limit. That is, samples with few stars cannot reveal extinction structures, while larger samples may detect weak jumps and additional extinctions from other components along the LoS. We establish an interval for $n$ to perform density adjustment, and samples outside of this interval will be fixed at a predefined lower or upper bound. Considering that a typical MC with an angular size $\rm\gtrsim 10-50~{}arcmin^{2}$ (Yuan et al., 2021; Zhang et al., 2024) can be measured within 3 kpc near the Galactic plane, we recommend a range of the number density of stars as $n\in\left(0.2,20\right){\rm pc}^{-1}$ and that the iteration stops when the number of stars in a subset is less than 20.
@@ 3.3 The Uncertainties of Distance by Bootstrap
Since the uncertainties of the distances of jump points cannot be directly derived from the data uncertainties, the bootstrap method offers a way to reflect systematic measurement errors in our results. A bootstrap is described by $X_{i}=\mu_{i}+\varepsilon$, where $X_{i}$ is the new dataset used for calculations, $\mu_{i}$ is the original value corresponding to distance or extinction in this study, and $\varepsilon$ represents the small perturbations.
@@
Here, we can primarily define an extinction jump as an increase in extinction values over a short distance interval. For instance, under typical distance uncertainties of 5%–10% ($\delta D\sim 50$ pc), $\Delta A_{V}$ reaches $\sim 0.15$ mag (see Section 5.1). The ratio between $\Delta A_{V}$ and $\delta D$ reflects a density of extinction, equivalent to $k\gtrsim 0.003~{}{\rm mag\cdot pc^{-1}}$ (see Appendix A.3). Obviously, the definition of jumps is mainly limited by the uncertainties of star samples. Fortunately, in real data, the increase in extinction ($\sim 0.75~{}{\rm mag\cdot kpc^{-1}}$; Lynga, 1982; Girardi et al., 2005) brought by the slope $k$ is generally less than the extinction feature caused by jumps.
Figure 2: An illustration of our multi-extinction-jump model (here in Mode I, $C$=3 for the left column, $C$=1 for the middle and right columns) applied to the mock data. Different rows set the jump points number from 1 to 6, with distances at $\left[300,700,1200,1800,2500,3300\right]{\rm pc}$, respectively. The three columns mirror different LoS extinction properties (see the text). Grey dots are the produced mock data, and black dots are binned values every 20 pc.
@@ 4.1 Validation Using Mock Data
To verify the ability of our method on detecting weak jumps, we generate mock data to simulate the distribution of stars in the $D$-$A$ diagram. Firstly, we categorize the synthetic extinctions into different distance intervals with distinct mean and dispersion. Both the distribution of distance errors and $A_{G}$ errors follow a Truncated Gaussian model. Referring to stars in Cygnus (from Gaia DR3 in Section 2.2.1), we found that the distance error is linearly correlated with distance, with a coefficient of approximately 0.1 and a dispersion of about 0.06. As for $A_{G}$ errors, their average and dispersion show no apparent correlation with $A_{G}$ itself. We have therefore set them to 0.04 mag for both the average and the dispersion based on the data toward the Cygnus.
@@
In Figure 2, the number of jump points increases from one to six in different rows from top to bottom. For the columns, the left column presents a simple multi-jump pattern, resembling a stairs with different star number densities $n=\left[0.2,0.3,0.4,0.5,0.4,0.3,0.2\right]{\rm pc}^{-1}$. The middle column gives smaller $\Delta A_{G}$ and larger $\sigma$ compared to left panels with lower number density $n=\left[0.1,0.15,0.2,0.25,0.2,0.15,0.1\right]{\rm pc}^{-1}$. The panels in right column are added a linear gradient with a slope $k\sim 3$$\times 10^{-5}$ $\rm mag\cdot pc^{-1}$ based on the middle column. The distances corresponding to each jump point are set to $\left[300,700,1200,1800,2500,3300\right]{\rm pc}$. Obviously, our model demonstrates good performance on these mock data (see the identified extinction jumps indicated by vertical lines in Figure 2).
Through statistical analysis of large number of mock data, we are able to quantitatively evaluate the optimized parameters of our method for various application scenarios. We set up three types of mock data (1, 2, and 3 jump points), with 300,000 groups of $D$-$A$ data sequences for each type. In addition to the uncertainties of distance and extinction setting before, we randomly combine the following settings for $D$, $\Delta A_{G}$, $\sigma$, $n$, and $k$ to obtain $D$-$A$ data sequences. These $D$-$A$ data sequences represent that the physical quantities are completely randomly generated and combined under a certain statistical distribution. For weak jumps with a linear extinction gradient along the LoS, we set up (1) stars’ distances in [10, 3000] pc; (2) jump points’ distances uniformly distributed in [60, 2950] pc; (3) $\Delta A_{G}\sim\xi(0.2)+0.15$ mag, where $\xi$ represent an exponential distribution; (4) $\sigma\sim\xi(0.2)+0.2$ mag; (5) $P(n|0.2,0.05)$ is Gaussian for $n$ in [0.1, 0.4] $\rm pc^{-1}$; (6) $P(k|3,1)$ is Gaussian for $k$ in [1, 6]×$\times 10^{-5}$ $\rm mag\cdot pc^{-1}$.
@@
By adjusting the parameter $C$, we analyze the three types of mock data to understand their behavior under various conditions. The detection of jump point in the mock data is categorized into the following three situations: detecting the preset jump points (True Positive; $TP$); failing to detect the preset jump points (False Negative; $FN$); and detecting non-preset change points (False Positive; $FP$). We determine whether a jump point has been successfully detected by seeing if the difference from the preset value is within the $1\sigma$ confidence interval. The reliability ($precision$) of detection is the proportion of $TP$ among all detected change points $presicion=\frac{TP}{TP+FP}$, and the completeness ($recall$) is the proportion of $TP$ among all preset jump points $recall=\frac{TP}{TP+FN}$. The $F1$ score, which serves as a metric for the model’s accuracy in classification, can be derived from $precision$ and $recall$ $F_{1}=2\times\frac{precision\times recall}{precision+recall}$.
@@
Table 1 lists the mean of all the statistical results, the high $precision$, $recall$, and $F1$ score presents good performance for detecting weak jump points with slopes for $C=1$. Raising $C$ will improve the $precision$, but the $recall$ decreases significantly. In addition, the mean $F1$ score of the first detected jump point shows no significant difference compared to the scores of all detections. This indicates that early-detected jump points also exhibit robustness, while the number of jump points has negligible impact on the jump point detection, even though the downward trend of $F1$ score with an increasing number of jump points reveals some issues (see Appendix A.1). The hypothesis testing with mock data indicates that controlling a single parameter ($C$) helps this method to detect fewer erroneous change points.
@@
$C$ $\rm Type^{a}$ $TP$ $FP$ $FN$ $Precision$ $Recall$ $F1$ score $\rm Gap^{b}$ $F1$ $\rm score_{I}^{c}$
@@ 4.2 Validation Using Distances of Masers
The dense molecular clumps associated with masers can cause significant stellar extinction. When applying the CUSUM-based model to stars along the LoS toward masers, we expect to detect jumps in $D$-$A$ likely caused by the associated MCs. We use maser samples from Reid et al. (2019) and Hyland et al. (2023) to correspond the distance of HMSFRs. In total, over 200 of masers have had their parallaxes and proper motions measured. Considering the precision of Gaia’s geometric measurements of stars, we selected 75 maser samples with parallaxes greater than 0.333 mas for validation.
We selected all stars with a relative parallax error smaller than 0.2, centered on the locus of each maser within a projection radius of $30^{\circ}.$3. The maximum distance cutoff was set at 5 kpc. We applied the Mode I method, which involves setting a universal constant $C=3$. Here the scaling factor $C=3$ is quite strict for identifying the jumps caused by the majority of these HMSFRs traced by maser measurements, with $n$ in the range of $\left(0.2,0.6\right){\rm pc}^{-1}$. We manually match the distances derived from maser parallaxes with the locations where $A_{G}$ exhibited jumps. One may wonder whether the matching of jump points to maser distances is due to human factors. In the majority of our selected samples, there are only two or three jump points are observed within 5 kpc. Therefore, it is reasonable to consider the association between the identified jump points and masers with the corresponding HMSFRs is not coincidental (see examples in Figure C1). We conclude that our method can independently capture the dense gas structures with $\Delta A_{G}\gtrsim 0.5$ mag based on the well cross-match between the extinction distances and masers parallax distances.
@@
Finally, 46/75 (42/58) of masers within 3 kpc (2.5 kpc) are found to correspond with the jump points identified by our Mode I. A good linear relationship between all the detected distances and those derived from masers are well presented in Figure 3 without systematic deviation. Most of the differences between two methods fall within the statistical tolerance of a $1\sigma$ error. Among them, for some samples with small number density, we lowered the scaling factor appropriately and successfully recovered another 6 samples (5 cases with $C=2$ and 1 with $C=1$). The consistency in the figure indicates that our method can well trace the distance of the dense regions of HMSFRs.
Our method fails to detect approximately 28% of the sample within 2.5 kpc and 76% within 2.5–3 kpc range. This is attributed to a significant decrease in the completeness of stars with precise parallax and $A_{G}$ measurements, particularly due to the “extinction wall” phenomenon (Straizys et al., 1993), where foreground dust obscures distant stars, especially those behind the Aquila Rift and Cygnus Rift. Notwithstanding these limits, in the Cygnus region, our method successfully identified multiple extinction jumps associated with MCs linked to masers in star-forming regions beyond the Cygnus Rift (highlighted as orange points in Figure 3). This method proves more effective than the Bayesian single-extinction-jump method (Zhang et al., 2024) in detecting extinction components at greater distances with fewer background stars. Furthermore, as distance increases, especially beyond 2.5 kpc, the inability to detect the more distant extinction jump points is not due to mismatches with maser distance measurements, but rather because of insufficient stellar sample density in these regions.

### v2

@@ 3.2 Detection of Weak Jumps in D-A Plot
Considering the end of the iterative procedure in Steps 3 and 4 (as an example of Figure 1f), for a set of $D$-AA data with a single jump point, we can obtain $S_{\rm diff}^{0}$ according to the definition
@@
$\xi\left(\Delta A_{\lambda},\sigma\right)$ exhibits a complex analytic form. However, $\xi\left(\Delta A_{\lambda},\sigma\right)\propto\sigma$ when $\sigma\gtrsim\Delta A_{\lambda}$. We focus on analyzing two primary cases. When $\Delta A_{\lambda}\gg\sigma$, we define these features as significant extinction-jump points. The algorithm facilitates efficient traversal of all jump points on the $D$-AA diagram (see examples in Figure 2). For the other case, due to intrinsic dispersion and uncertainty, the extinction of stars in Gaia data leads to a magnitude uncertainty of $\sigma$ that is comparable to most of the $\Delta A_{\lambda}$ values we can detect as jump points. In observations, these jump points are collectively referred to as “weak jumps”, and statistically the number of these weak jumps accounts for the majority. Additionally, the $D$-AA diagram includes gentle slopes ($k$) that may be caused by diffuse and extended gas along the LoS, which exhibit more linear properties in $D$-AA space and present challenges for the CU-JADE method (see Appendix A.3). In the following subsections, we will address the influence of these factors on the results and optimize the parameter settings to identify these extinction features with weak jumps.
@@ 3.2.1 Comparable $\Delta A_{\lambda}$ and $\sigma$ for Weak Jumps
By Equations (6), (A7) and Figure A1, when $\Delta A_{\lambda}$ and $\sigma$ are the same magnitude, $\sigma$ has a major influence on the $S_{\rm diff}$. Approximately, we have $S_{\rm diff}~\propto~\sqrt{n}~\sigma,~\sigma\gtrsim\Delta A_{\lambda}$. The influence of $\Delta A_{\lambda}$ on $S_{\rm diff}$ is negligible compared to $S_{\rm diff}^{0}$. Combining Equation (5), we have
@@
The $CL$ for an extinction-jump is proportional to the jump value and the square root of number density, but inversely proportional to $A_{\lambda}$ dispersion. Therefore, a larger sample size tends to make it easier to detect a jump point, as well as the larger $\Delta A_{\lambda}$. A large dispersion would smear out the jump when it is comparable to the $\Delta A_{\lambda}$ values. Since $\Delta A_{\lambda}$ and $\sigma$ depend upon the LoS extinctions and the uncertainties of data, the value of $CL$ now is only influenced by the number density of the samples in the calculation.
@@ 3.2.2 Two Modes for Stellar Sample Density Variation
The second mode (Mode II) adjusts the $CL$ threshold with the number density, scaling with $\sqrt{n}$, to mitigate the influence of star density. In Mode II, we refine our method to apply varying $CL$ thresholds in each split iteration, enabling the jump point detection in intervals with sparser star distribution. In cases with prior knowledge of the target area, we recommend the Mode II. It is important to control bias stemming from sample density. The number density $n$ is calculated based on the star samples number per unit length in a specific region and distance interval, $n=\frac{N}{d_{max}-d_{min}}$, where $N$ is the number of stars, $d_{max}$ and $d_{min}$ are the maximum and minimum distances of selected stars in each interval, respectively.
As discussed in Section 3.1, to enhance the $CL$ while mitigating false detections, we multiply $S_{\rm diff}$ by a factor $C$ ($C\sqrt{\beta n}$ for Mode II, $\beta$ is a constant for fine-tuning). $C$ is usually set to be 131-3, considering $\Delta A_{\lambda}$ ranges from $\left[0.5,5\right]$ mag and $n$ ranges from $\left[0.1,10\right]~{\rm pc}^{-1}$. In the cases of large dispersion $\sigma$, or too small $\Delta A_{\lambda}$ due to weak extinctions, we have to do coarse tuning to the criteria by reducing the value of $C$, as $CL$ decreases rapidly during the iterations. We find that $C=1$ is more effective for weak jumps (see details in Section 4.1). Fine-tuning of the threshold can also be achieved by altering the percentile value of $S_{\rm diff}^{0}$ in shuffled $S_{\rm diff}$ (generally taken as 0.95). Mode II has to be complemented with both a lower and an upper limit. That is, samples with few stars cannot reveal extinction structures, while larger samples may detect weak jumps and additional extinctions from other components along the LoS. We establish an interval for $n$ to perform density adjustment, and samples outside of this interval will be fixed at a predefined lower or upper bound. Considering that a typical MC with an angular size $\rm\gtrsim 10-50~arcmin^{2}$ (Yuan et al., 2021; Zhang et al., 2024) can be measured within 3 kpc near the Galactic plane, we recommend a range of the number density of stars as $n\in\left(0.2,20\right){\rm pc}^{-1}$ and that the iteration stops when the number of stars in a subset is less than 20.
@@ 3.3 The Uncertainties of Distance by Bootstrap
Since the uncertainties of the distances of the jump points cannot be directly derived from the data uncertainties, the bootstrap method offers a way to reflect systematic measurement errors in our results. A bootstrap is described by $X_{i}=\mu_{i}+\varepsilon$, where $X_{i}$ is the new dataset used for calculations, $\mu_{i}$ is the original value corresponding to distance or extinction in this study, and $\varepsilon$ represents the small perturbations.
@@
Here, we can primarily define an extinction-jump as an increase in extinction values over a short distance interval. For instance, under typical distance uncertainties of 5%–10% ($\delta D\sim 50$ pc), $\Delta A_{V}$ reaches $\sim 0.15$ mag (see Section 5.1). The ratio between $\Delta A_{V}$ and $\delta D$ reflects a density of extinction, equivalent to $k\gtrsim 0.003~{\rm mag\cdot pc^{-1}}$ (see Appendix A.3). Obviously, the definition of jumps is mainly limited by the uncertainties of star samples. Fortunately, in real data, the increase in extinction ($\sim 0.75~{\rm mag\cdot kpc^{-1}}$; Lynga, 1982; Girardi et al., 2005) brought by the slope $k$ is generally less than the extinction feature caused by jumps.
Figure 2: An illustration of our multi-extinction-jump model (here in Mode I, CC=3 for the left column, CC=1 for the middle and right columns) applied to the mock data. Different rows set the jump points number from 1 to 6, with distances at $\left[300,700,1200,1800,2500,3300\right]{\rm pc}$, respectively. The three columns mirror different LoS extinction properties (see the text). Gray dots are the produced mock data and black dots are binned values every 20 pc.
@@ 4.1 Validation Using Mock Data
To verify the ability of our method to detect weak jumps, we generate mock data to simulate the distribution of stars in the $D$-AA diagram. Firstly, we categorize the synthetic extinctions into different distance intervals with distinct mean and dispersion. Both the distribution of distance errors and $A_{G}$ errors follow a truncated Gaussian model. Referring to stars in Cygnus (from Gaia DR3 in Section 2.2.1), we found that the distance error is linearly correlated with distance, with a coefficient of approximately 0.1 and a dispersion of about 0.06. As for $A_{G}$ errors, their average and dispersion show no apparent correlation with $A_{G}$ itself. We have therefore set them to 0.04 mag for both the average and the dispersion based on the data toward the Cygnus.
@@
In Figure 2, the number of jump points increases from one to six in different rows from top to bottom. For the columns, the left column presents a simple multijump pattern, resembling a stairs with different star number densities $n=\left[0.2,0.3,0.4,0.5,0.4,0.3,0.2\right]{\rm pc}^{-1}$. The middle column gives smaller $\Delta A_{G}$ and larger $\sigma$ compared to left panels with lower number density $n=\left[0.1,0.15,0.2,0.25,0.2,0.15,0.1\right]{\rm pc}^{-1}$. The panels in right column have an added a linear gradient with a slope $k\sim 3$ $\times 10^{-5}$ $\rm mag\cdot pc^{-1}$ based on the middle column. The distances corresponding to each jump point are set to $\left[300,700,1200,1800,2500,3300\right]{\rm pc}$. Obviously, our model demonstrates good performance on these mock data (see the identified extinction jumps indicated by vertical lines in Figure 2).
[…]
@@
By adjusting the parameter $C$, we analyze the three types of mock data to understand their behavior under various conditions. The detection of jump point in the mock data is categorized into the following three situations: detecting the preset jump points (True Positive; $TP$); failing to detect the preset jump points (False Negative; $FN$); and detecting non-preset change points (False Positive; $FP$). We determine whether a jump point has been successfully detected by seeing if the difference from the preset value is within the $1\sigma$ confidence interval. The reliability (precisionprecision) of detection is the proportion of $TP$ among all detected change points $presicion=\frac{TP}{TP+FP}$, and the completeness (recallrecall) is the proportion of $TP$ among all preset jump points $recall=\frac{TP}{TP+FN}$. The $F1$ score, which serves as a metric for the model’s accuracy in classification, can be derived from precisionprecision and recallrecall $F_{1}=2\times\frac{precision\times recall}{precision+recall}$.
@@
Table 1 lists the mean of all the statistical results, the high precisionprecision, recallrecall, and $F1$ score presents good performance for detecting weak jump points with slopes for $C=1$. Raising $C$ will improve the precisionprecision, but the recallrecall decreases significantly. In addition, the mean $F1$ score of the first detected jump point shows no significant difference compared to the scores of all detections. This indicates that early-detected jump points also exhibit robustness, while the number of jump points has negligible impact on the jump point detection, even though the downward trend of $F1$ score with an increasing number of jump points reveals some issues (see Appendix A.1). The hypothesis testing with mock data indicates that controlling a single parameter ($C$) helps this method to detect fewer erroneous change points.
@@
$C$ $\rm Type^{a}$ TPTP FPFP FNFN PrecisionPrecision RecallRecall $F1$ score $\rm Gap^{b}$ $F1$ $\rm score_{I}^{c}$
@@ 4.2 Validation Using Distances of Masers
The dense molecular clumps associated with masers can cause significant stellar extinction. When applying the CUSUM-based model to stars along the LoS toward masers, we expect to detect jumps in $D$-AA that are likely caused by the associated MCs. We use maser samples from Reid et al. (2019) and Hyland et al. (2023) to correspond the distance of HMSFRs. In total, over 200 of masers have had their parallaxes and proper motions measured. Considering the precision of Gaia’s geometric measurements of stars, we selected 75 maser samples with parallaxes greater than 0.333 mas for validation.
We selected all stars with a relative parallax error smaller than 0.2, centered on the locus of each maser within a projection radius of $0^{\circ}.3$. The maximum distance cutoff was set at 5 kpc. We applied the Mode I method, which involves setting a universal constant $C=3$. Here, the scaling factor $C=3$ is quite strict for identifying the jumps caused by the majority of these HMSFRs traced by maser measurements, with $n$ in the range of $\left(0.2,0.6\right){\rm pc}^{-1}$. We manually match the distances derived from maser parallaxes with the locations where $A_{G}$ exhibited jumps. One may wonder whether the matching of jump points to maser distances is due to human factors. In the majority of our selected samples, there are only two or three jump points are observed within 5 kpc. Therefore, it is reasonable to consider that the association between the identified jump points and masers with the corresponding HMSFRs is not coincidental (see examples in Figure C1). We conclude that our method can independently capture the dense gas structures with $\Delta A_{G}\gtrsim 0.5$ mag based on the well crossmatch between the extinction distances and masers parallax distances.
@@
Finally, 46/75 (42/58) of masers within 3 kpc (2.5 kpc) are found to correspond with the jump points identified by our Mode I. A good linear relationship between all the detected distances and those derived from masers are well presented in Figure 3 without systematic deviation. Most of the differences between two methods fall within the statistical tolerance of a $1\sigma$ error. Among them, for some samples with small number density, we lowered the scaling factor appropriately and successfully recovered another six samples (five cases with $C=2$ and one with $C=1$). The consistency in the figure indicates that our method can well trace the distance of the dense regions of HMSFRs.
Our method fails to detect approximately 28% of the sample within 2.5 kpc and 76% within 2.5–3 kpc range. This is attributed to a significant decrease in the completeness of stars with precise parallax and $A_{G}$ measurements, particularly due to the “extinction wall” phenomenon (Straizys et al., 1993), where foreground dust obscures distant stars, especially those behind the Aquila Rift and Cygnus Rift. Notwithstanding these limits, in the Cygnus region, our method successfully identified multiple extinction jumps associated with MCs linked to masers in star-forming regions beyond the Cygnus Rift (highlighted as orange points in Figure 3). This method proves more effective than the Bayesian single-extinction-jump method (Zhang et al., 2024) in detecting extinction components at greater distances with fewer background stars. Furthermore, as distance increases, especially beyond 2.5 kpc, the inability to detect the more distant extinction-jump points is not due to mismatches with maser distance measurements but rather because of insufficient stellar sample density in these regions.